OpenStudy (anonymous):

(a^-1)-(b^-1) ------------ (b^-2)-(a^-2)

OpenStudy (anonymous):

the line is divide

OpenStudy (anonymous):

\[\frac{\frac{1}{a}-\frac{1}{b}}{\frac{1}{b^2}-\frac{1}{a^2}}\] multiply top and bottom by \[a^2b^2\] get \[\frac{ab^2-a^2b}{a^2-b^2}\] \[\frac{ab(b-a)}{(a-b)(a+b)}\] and finally \[-\frac{ab}{a+b}\] since \[\frac{b-a}{a-b}=-1\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

So the answer is -1? Also how did you get from the second to last step to the last step?